What is the positive difference of the root of x2-7x-9=0

Grade:12th pass

1 Answers

Aastha Singh
3 years ago
Quadratic equation is
x2−7x−9.
a+β=7,aβ=−9
(a−β)2=(a+β)2−4aβ
=49+36=85
hence , a−β=85​